# Reinforcement learning for anisotropic p-adaptation and error estimation in high-order solvers

### v1

@@ Abstract
We present a novel approach to automate and optimize anisotropic p-adaptation in high-order h/p solvers using Reinforcement Learning (RL). The dynamic RL adaptation uses the evolving solution to adjust the high-order polynomials. We develop an offline training approach, decoupled from the main solver, which shows minimal overcost when performing simulations. In addition, we derive a RL-based error estimation approach that enables the quantification of local discretization errors. The proposed methodology is agnostic to both the computational mesh and the partial differential equation being solved.
The application of RL to mesh adaptation offers several benefits. It enables automated, adaptive mesh refinement, reducing the need for manual intervention. It optimizes computational resources by dynamically allocating high-order polynomials where necessary and minimizing refinement in stable regions. This leads to computational cost savings while maintaining solution accuracy. Furthermore, RL allows for the exploration of unconventional mesh adaptations, potentially enhancing the accuracy and robustness of simulations. This work extends our original research in [1], offering a more robust, reproducible, and generalizable approach applicable to complex three-dimensional problems. We provide validation for laminar and turbulent cases: circular cylinders, Taylor Green Vortex and a 10MW wind turbine to illustrate the flexibility of the proposed approach.
@@ 1 Introduction
When considering flow simulations, where unsteady and turbulent flows evolve in time and space, RL shows great potential, since it offers the possibility to automatically select numerical parameters and mesh resolutions as flows evolve. An example is the selection of constants in turbulence models [8, 9], optimal parameters in high-order schemes [10, 11] or mesh adaptation.
@@
Mesh adaptation or adaptive mesh refinement (AMR) is essential in numerical methods, allowing for the refinement or coarsening of the computational mesh based on solution smoothness and computational cost. Traditionally, this process has depended on heuristics or manual intervention, which can be time-consuming and limited in capturing complex flow phenomena. The integration of RL for mesh adaptation provides a way of automating this process, but it is still in its early stages. RL has been successfully used for mesh adaptation (h-refinement) in recent researches, where agents were trained to increase the number of elements in the mesh, increasing the accuracy in critical regions. In particular, Foucart et al. [12] compared deep reinforcement learning with heuristic approaches to perform an h-adaptive mesh refinement that is generalizable for different partial differential equations (PDE). Also, in [13] new policy architectures are considered, which are independent of the mesh size and are trained directly from simulations. In the field of multi-agent RL, Dzanic et al. [14] show a novel approach based on independent PPO that can be used for h- and p-refinement, with the limitation that only periodic meshes are used and a single level of refinement is allowed. These approaches lead to accurate solutions with a reduced computational cost compared to traditional manual h-mesh refinement.
When dealing with high-order solvers, utilizing RL for p-mesh adaptation brings about benefits. RL-driven p-adaptation optimizes computational resources by intelligently allocating high-order polynomials where required and minimizing refinement in regions with stable flow. This strategic allocation of resources leads to significant computational cost savings while maintaining the precision of the solution.
@@
This work builds on our foundational research presented in [1], where we established the basis for applying reinforcement learning to p-adaptation in 1D partial differential equations. We have further refined the original concept, resulting in a more robust and reproducible approach that can tackle complex three-dimensional problems. Here, we focus on the Navier-Stokes equations, but the methodology is flexible and could be used to solve other PDEs. We explore the uncharted domain of anisotropic high-order p-adaptation using RL, with the aim of automating and optimizing the mesh adaptation process. By incorporating RL algorithms into simulations, we can train agents to adjust the polynomial order in high-order solvers, according to evolving solutions. The RL agent interacts with the numerical solver, observes the flow state, and makes decisions on mesh modifications. During offline training, the agent receives rewards based on the accuracy and efficiency of the numerical solution. This training is decoupled from the real simulation (the solver is not required during the training), enhancing the efficiency of the proposed approach. Furthermore, this novel methodology allows us to train highly generalizable RL agents that are mesh independent and can be used to potentially solve any PDE with a high-order solver.
@@
Finally, to contextualize the novelty of our work, we summarize in Table 1 the main characteristics of our RL based p-adaptation strategy and other state-of-the-art RL-based adaptive mesh refinement strategies proposed in [12] and [14]. It can be seen that our work considers p-adaptation and is trained in 1D problems but can be used for 3D flows, resulting in computational cost savings during the training. We do not need to retrain the RL agent for new cases; once it is trained, it works for any (tested) case. In addition, we propose using a value iteration algorithm that is fully reproducible and robust (since it does not rely on neural networks). Finally, we provide an RL-based error estimation, and apply the p-adaptation to 3D turbulent cases (not done previously).
@@
The rest of the paper is organized as follows: first, in Section 2 we comment the state-of-the-art in p-adaptation for high-order solvers. In Sections 3.1 and 3.2 we provide some background regarding discontinuous Galerkin solvers and the reinforcement learning framework. Then, in Section 3.3 we describe the main RL algorithm that is the core of the whole methodology, and in Section 3.4 we define all the required elements that are necessary to compute the previous algorithm. In Section 3.5, all the underlying information is provided in relation to the implementation of the previous methodology in a real solver. Finally, in Section 4 we show different results to validate the proposed strategy and in Section 5 we provide some final conclusions.
@@ 2 State of the art in p-adaptation
Adaptation strategies in computational fluid dynamics are typically categorized based on the type of error measurement that is employed. These categories include feature-based adaptation, adjoint-based adaptation, and local error-based adaptation. Comparative studies of these approaches have been conducted by Fraysse et al. [15] for finite-volume approximations and by Kompenhans et al. [16] and Naddei et al. [17] for high-order discontinuous Galerkin (DG) methods.
@@
The adjoint-based adaptation defines a functional target (e.g. drag or lift in external flow aerodynamics). The adjoint problem is then solved to obtain a spatial distribution of the functional error, which guides the mesh adaptation. This technique, originally developed for structural analysis using the Finite Element Method (FEM) by Babuška and Miller [24, 25], has been applied to adaptation strategies in DG methods [26, 27, 28]. Despite its sophistication, this approach has the drawback of high computational cost and storage requirements to solve the adjoint problem and store error estimators, especially in unsteady flows. Moreover, it only guarantees the reduction of the analyzed functional error, which could deteriorate the error of other functionals.
@@
Other techniques used to dynamically adapt the dissipation properties of high-order methods include troubled cell-indicators, multidimensional optimal order detection (MOOD), and adaptive dissipation control. Troubled cell-indicators were initially introduced in [35] and are commonly used in conjunction with TVD-TVB limiters [36, 21, 37]. MOOD methods assess the maximum achievable accuracy order at each cell that ensures the fulfillment of physical and numerical admissibility conditions [38, 39, 40, 41, 42, 43]. Adaptive dissipation control, on the other hand, aims at adapting the dissipation properties of the scheme by means of modal filters (which are closely related to p-adaptation) [44, 45, 46, 47].
@@ 3.1 High-order discontinuous Galerkin solvers
In this work we use the nodal discontinuous Galerkin variant called DGSEM (Discontinuous Galerkin Spectral Element Method) [48]. DG methods are characterized by their ability to use mesh refinement through polynomial enrichment (p-adaptation) to achieve highly accurate solutions; that is, low numerical errors. Such high-order polynomial methods produce an exponential decay of the error for sufficiently smooth solutions instead of the algebraic decay characteristic of low-order techniques [49]. The DGSEM method allows us to compute the approximated solution in each element of the computational mesh, which is a tessellation of non-overlapping hexahedral elements for a 3D case, through Lagrange polynomials. Within this framework, the elements of the mesh have to be transformed from their original shape to a cube, which is used as a reference element $E=[-1,1]^{3}$, as shown graphically in Figure 1. Once the equations are solved in computational space, the inverse transform is applied to recover the original geometry of the problem. The geometric transformation from the physical space to the computational space provides a mechanism to use curvilinear meshes up to an arbitrary order. We provide more details on the formulation in A and B.
@@ 3.2 Reinforcement learning framework
    Agent: The brain behind the reinforcement learning which decides the best action to take. The agent is modeled with the policy function $\pi(s)$, which is obtained through the training process and determines the action $a$ to take from the state $s$.
@@
While the definitions of the environment and the actions are straightforward, the state and the reward have to be cautiously designed to obtain good performance. These elements of the RL framework are explained in detail in Sections 3.4.1, 3.4.2, and 3.4.3.
@@ 3.3 The value iteration algorithm
In this research, we use a family of model-based algorithms known as Dynamic Programming (DP) to solve the RL problem. These algorithms allow to compute optimal policies given a known model of the environment as a finite MDP [4]. Therefore, the probability transition function $\mathbb{P}(s^{\prime},r\mid s,a)$ for the environment is explicitly required. DP does not rely on neural networks (NN), but uses tabular representation of the variables of interest instead, which limits the size of the state space. This limitation has driven the development of Deep Reinforcement Learning methods. However, classic reinforcement learning methods like DP excel in robustness, reproducibility, and reliability, along with fast convergence, all of which are highly desirable for numerical simulations. Therefore, whenever feasible, these methods should be considered.
@@
The exact solution of the previous problem can be obtained by solving a system of $N$ nonlinear equations (due to the nonlinearity introduced by the $\max$ operator), with $N$ as the total number of states in $\mathcal{S}$. In addition, there are different approaches to solving the problem approximately using an iterative algorithm. Among the available options, the value iteration algorithm provides an update rule for $v(s)$, based on the Bellman equation, that converges to the optimum $v^{*}$ for an arbitrary initialization $v_{0}$ [4]:
@@ 3.4 Reinforcement learning strategy for p-adaptation
The main objective of this research is to provide a flexible p-adaptation RL agent that can be used to solve a wide variety of problems (or PDE) and that is not dependent on the computational mesh. Furthermore, the resulting agent, once trained, should be used as it is in any high-order solver and it should provide accurate results for every possible scenario. This approach would require training the agent in a way that is common for every problem, which is not a simple task. In this work, we have defined the state and the reward to provide enough information to the agent, while being general enough. In this way, the same agent can be used for different problems.
@@ 3.4.1 State definition
As our goal is to create an agent as general as possible, the state must be as general as possible. Here, we consider a nodal-based discontinuous Galerkin solver, as was explained in Section 3.1. Therefore, we have to focus on the only thing that is common for every DGSEM solver: the solution is computed at the Gauss nodes inside each element. Hence, the state is simply defined as the values of the variable of interest, $q$, at the Gauss nodes, as represented in Figure 5(a). The size of the resulting state is $p+1$ (the same as the number of Gauss nodes) and is different depending on the polynomial order. This approach considers a different state in each element of the computational mesh; that is, the RL agent will handle each element independently, leading to a mesh-independent strategy. However, the main issue behind this definition of the state is that, in a general problem, $q$ may take any value, since $q\in\mathbb{R}$. This would lead to an unbounded and continuous state space, which does not match neither the MDP framework nor the value iteration algorithm, as they both deal with a finite state space. Hence, the possible values of the variable of interest $q$ must be discretized, so we can define an appropriate finite space state $\mathcal{S}$.
The proposed methodology takes the values of $q$ at every node within an element and normalizes the values from $-1$ to $1$. In this way, for each element there will always be one node with a value of $-1$ and another with a value of $1$. If it is the case that the value of $q$ is exactly the same at every node (below a very small tolerance, then the value at each node is set to 00 (see Section 3.5 for additional implementation details). Then, the state is discretized in $N_{l}$ levels. The higher the value of $N_{l}$, the more accurate the final agent will be, but the slower the training phase. After some trial and error, we discovered that a value of $N_{l}=11$ provides a good trade-off.
@@ 3.4.3 Reward definition
The definition of the reward will be influenced by the framework used to implement the p-adaptation strategy. In general, the p-adaptation process in a high-order solver is highly time-consuming. Therefore, we cannot afford to adapt the mesh every iteration. Furthermore, for numerical stability reasons, each time the mesh is adapted, only one action is taken in each element; that is, the polynomial order can be increased or decreased by 1 unit (see Section 3.5 for additional details regarding the implementation). These details should be considered when defining an appropriate definition of the reward function.
Additionally, we want the agent to find the optimal polynomial order for each element of the computational mesh, but it is not relevant how it reaches that polynomial, and therefore the reward must not depend on the action. However, it is important to decide whether the reward should depend on the current state $s$, or on the next state $s^{\prime}$. Both approaches are possible and valid, but the meaning is slightly different in each case. The key behind this decision is based on the knowledge that the trained agent will rely on the value function $v(s)$ to follow the optimal policy, which is simply to select the action that leads to a state with a higher $v(s^{\prime})$ (see the previous Section 3.3). Furthermore, as it is stated in eq. (2), each value function is directly linked to the reward $r(s,a,s^{\prime})$, and hence there is a direct correlation between the selected reward and the final policy.
@@
Within this framework, if we choose the second reward option, $r=r(s^{\prime})$, the agent will try to find a state from which it is possible to reach a highly beneficial future state. However, if we can only perform one action at a time, the agent will not be able to reach the desired state until we can adapt the second time, at least. By the time we adapt again (several iterations later), the flow field might have changed, and that desirable future state might not be beneficial anymore. In contrast, if we define the reward as $r=r(s)$, the agent will take the action that leads to a highly beneficial state right now, even if the future is uncertain. In this way, at least, we can ensure that the chosen action is the best possible one right now. This is the reason why we have selected this approach when defining the reward function.
@@
where $\sigma$ is the standard deviation and $\alpha$ is a control parameter. The most important contribution to the reward function is the exponential part, which measures the accuracy (or the error). This part will drop the reward to zero if the $rmse$ is big compared to the standard deviation, $\sigma$. The second contribution is the cost, and it will only be relevant when the $rmse$ is small enough. This part will decrease the reward when a high polynomial order is used, as it will require a high computational cost. Therefore, the reward function prioritizes the reduction of the error; but if the error is acceptable (based on the value of $\sigma$), then the polynomial degree is reduced as much as possible (without significantly increasing the error).
On the one hand, the value of the parameter $\alpha$ provides a way to measure the trade-off between accuracy and computational cost. When $\alpha$ is small, the accuracy is more important than cost and vice versa. On the other hand, the value of the parameter $\sigma$ provides an estimate of the error that must be achieved in one element before the cost is taken into account. Therefore, the smaller the value of $\sigma$ the more important is the accuracy over the cost. Based on preliminary tests (not shown here), we have selected $\alpha=0.9$ and $\sigma=0.05$, which provides an appropriate balance between efficiency and accuracy of the p-adaptation strategy.

### v2

@@ Abstract
We present a novel approach to automate and optimize anisotropic p-adaptation in high-order h/p solvers using Reinforcement Learning (RL). The dynamic RL adaptation uses the evolving solution to adjust the high-order polynomials. We develop an offline training approach, decoupled from the main solver, which shows minimal overcost when performing simulations. In addition, we derive an inexpensive RL-based error estimation approach that enables the quantification of local discretization errors. The proposed methodology is agnostic to both the computational mesh and the partial differential equation to be solved.
The application of RL to mesh adaptation offers several benefits. It enables automated and adaptive mesh refinement, reducing the need for manual intervention. It optimizes computational resources by dynamically allocating high-order polynomials where necessary and minimizing refinement in stable regions. This leads to computational cost savings while maintaining the accuracy of the solution. Furthermore, RL allows for the exploration of unconventional mesh adaptations, potentially enhancing the accuracy and robustness of simulations. This work extends our original research in [1], offering a more robust, reproducible, and generalizable approach applicable to complex three-dimensional problems. We provide validation for laminar and turbulent cases: circular cylinders, Taylor Green Vortex and a 10MW wind turbine to illustrate the flexibility of the proposed approach.
@@ 1 Introduction
When considering flow simulations, where unsteady and turbulent flows evolve in time and space, RL shows great potential, since it offers the possibility of automatically selecting numerical parameters and mesh resolutions as flows evolve. An example is the selection of constants in turbulence models [8, 9], optimal parameters in high-order schemes [10, 11] or mesh adaptation.
@@
Mesh adaptation or adaptive mesh refinement (AMR) is essential in numerical methods, allowing the refinement or coarsening of the computational mesh based on solution smoothness and computational cost. Traditionally, this process has depended on heuristics, manual intervention or error estimators, which can be time-consuming and limited in capturing complex flow phenomena. A review of the current state of the art is provided in Section 2. The integration of RL for mesh adaptation offers a novel approach to automate the process, minimizing user intervention while providing a flexible and general framework applicable to a wide range of problems. However, it is still in its early stages of development. RL has been successfully used for mesh adaptation (h-refinement) in recent works, where agents were trained to increase the number of elements in the mesh, increasing the accuracy in critical regions. In particular, Foucart et al. [12] compared deep reinforcement learning with heuristic approaches to perform an h-adaptive mesh refinement that is generalizable for different partial differential equations (PDE). In addition, in the research conducted by Yang et al. [13] new policy architectures are considered, which are independent of the mesh size and are trained directly from simulations. In the field of multi-agent RL, Dzanic et al. [14] showed a novel approach based on independent PPO that can be used for h- and p-refinement, with the limitation that only periodic meshes are used and a single level of refinement is allowed. These approaches lead to accurate solutions with reduced computational cost compared to traditional mesh refinement.
@@
This work builds on our foundational research presented in [1], where we established the basis for applying reinforcement learning to p-adaptation in 1D partial differential equations. Based on this original concept, we have developed a novel and improved methodology for RL-based p-adaptation, resulting in a more robust and reproducible approach that can tackle complex three-dimensional problems. Here, we focus on the Navier-Stokes equations, but the methodology is flexible and could be used to solve other PDEs. We explore the uncharted domain of anisotropic high-order p-adaptation using RL, with the aim of automating and optimizing the mesh adaptation process. By incorporating RL algorithms into simulations, we can train agents to adjust the polynomial order in high-order solvers, according to evolving solutions. The RL agent interacts with the numerical solver, observes the flow state, and makes decisions on mesh modifications. During offline training, the agent receives rewards based on the accuracy and efficiency of the numerical solution. This training is decoupled from the real simulation (the solver is not required during the training), enhancing the efficiency of the proposed approach. Furthermore, this novel methodology allows us to train highly generalizable RL agents that are mesh-independent and can be used to potentially solve any PDE with a high-order solver.
@@
Finally, to contextualize the novelty of our work, we summarize in Table 1 the main characteristics of our RL-based p-adaptation strategy and other state-of-the-art RL-based adaptive mesh refinement strategies proposed in [12] and [14]. It can be seen that our work considers p-adaptation and is trained in 1D problems but can be used for 3D flows, resulting in computational cost savings during the training. We do not need to retrain the RL agent for new cases; once trained, it works for any (tested) case. In addition, we propose using a value iteration algorithm that is fully reproducible and robust (since it does not rely on neural networks). Finally, we provide an inexpensive RL-based error estimation and apply the p-adaptation strategy to 3D turbulent cases, which has not been done previously.
@@
The rest of the paper is organized as follows: first, in Section 2 we comment the state-of-the-art in p-adaptation for high-order solvers. In Sections 3.1 and 3.2 we provide some background on discontinuous Galerkin solvers and the reinforcement learning framework. Then, in Section 3.3 we describe the main RL algorithm that is the core of the entire methodology, and in Section 3.4 we define all the required elements that are necessary to compute the previous algorithm. In Section 3.5, all the underlying information in relation to the implementation of the previous methodology in a real solver is provided. Finally, in Section 4 we show different results to validate the proposed strategy and in Section 5 we provide some final conclusions.
@@ 2 State of the art in p-adaptation
Adaptation strategies in computational fluid dynamics are typically categorized based on the type of error measurement that is used. These categories include feature-based adaptation, adjoint-based adaptation, and local error-based adaptation. Comparative studies of these approaches have been conducted by Fraysse et al. [15] for finite-volume approximations and by Kompenhans et al. [16] and Naddei et al. [17] for high-order discontinuous Galerkin (DG) methods.
@@
The adjoint-based adaptation defines a functional target (e.g., drag or lift in external flow aerodynamics). The adjoint problem is then solved to obtain a spatial distribution of the functional error, which guides the mesh adaptation. This technique, originally developed for structural analysis using the Finite Element Method (FEM) by Babuška and Miller [24, 25], has been applied to adaptation strategies in DG methods [26, 27, 28]. Despite its sophistication, this approach has the drawback of high computational cost and storage requirements to solve the adjoint problem and store error estimators, especially in unsteady flows. Moreover, it only guarantees the reduction of the analyzed functional error, which could deteriorate the error of other functionals.
@@
Other techniques used to dynamically adapt the dissipation properties of high-order methods include troubled cell indicators, multidimensional optimal order detection (MOOD), and adaptive dissipation control. Troubled cell indicators were initially introduced in [35] and are commonly used in conjunction with TVD-TVB limiters [36, 21, 37]. MOOD methods assess the maximum achievable accuracy order in each cell that ensures the fulfillment of physical and numerical admissibility conditions [38, 39, 40, 41, 42, 43]. Adaptive dissipation control, on the other hand, aims at adapting the dissipation properties of the scheme by means of modal filters (which are closely related to p-adaptation) [44, 45, 46, 47].
@@ 3.1 High-order discontinuous Galerkin solvers
In this work we use the nodal discontinuous Galerkin variant called DGSEM (Discontinuous Galerkin Spectral Element Method) [48]. DG methods are characterized by their ability to use mesh refinement through polynomial enrichment (p-adaptation) to achieve highly accurate solutions; that is, low numerical errors. Such high-order polynomial methods produce an exponential decay of the error for sufficiently smooth solutions instead of the algebraic decay characteristic of low-order techniques [49]. The DGSEM method allows us to compute the approximate solution in each element of the computational mesh, which is a tessellation of non-overlapping hexahedral elements for a 3D case, through Lagrange polynomials. Within this framework, the elements of the mesh have to be transformed from their original shape to a cube, which is used as a reference element $E=[-1,1]^{3}$, as shown graphically in Figure 1. Once the equations are solved in computational space, the inverse transform is applied to recover the original geometry of the problem. The geometric transformation from the physical space to the computational space provides a mechanism to use curvilinear meshes up to an arbitrary order. We provide more details on the formulation in A and B.
@@ 3.2 Reinforcement learning framework
    Agent: The brain behind reinforcement learning, which decides the best action to take. The agent is modeled with the policy function $\pi(s)$, which is obtained through the training process and determines the action $a$ to take from the state $s$.
@@
While the definitions of the environment and the actions are straightforward, the state and the reward must be carefully designed to obtain good performance. These elements of the RL framework are explained in detail in Sections 3.4.1, 3.4.2, and 3.4.3.
@@ 3.3 The value iteration algorithm
In this research, we use a family of model-based algorithms known as Dynamic Programming (DP) to solve the RL problem. These algorithms allow to compute optimal policies given a known model of the environment as a finite MDP [4]. Therefore, the probability transition function $\mathbb{P}(s^{\prime},r\mid s,a)$ for the environment is explicitly required. DP does not rely on neural networks (NN), but uses tabular representation of the variables of interest instead, which limits the size of the state space. This limitation has driven the development of Deep Reinforcement Learning methods. However, classic reinforcement learning methods such as DP excel in robustness, reproducibility, and reliability, along with fast convergence, all of which are highly desirable for numerical simulations. Therefore, whenever feasible, these methods should be considered.
@@
The exact solution to the previous problem can be obtained by solving a system of $N$ nonlinear equations (due to the nonlinearity introduced by the $\max$ operator), with $N$ as the total number of states in $\mathcal{S}$. In addition, there are different approaches to solving the problem approximately using an iterative algorithm. Among the available options, the value iteration algorithm provides an update rule for $v(s)$, based on the Bellman equation, that converges to the optimum $v^{*}$ for an arbitrary initialization $v_{0}$ [4]:
@@ 3.4 Reinforcement learning strategy for p-adaptation
The main objective of this research is to provide a flexible p-adaptation RL agent that can be used to solve a wide variety of problems (or PDE) and that is not dependent on the computational mesh. Furthermore, the resulting agent, once trained, should be used as it is in any high-order solver, and it should provide accurate results for every possible scenario. This approach would require training the agent in a way that is common for every problem, which is not a simple task. In this work, we have defined the state and the reward to provide enough information to the agent, while being general enough. In this way, the same agent can be used for different problems.
@@ 3.4.1 State definition
As our goal is to create an agent as general as possible, the state must be as general as possible as well. Here, we consider a nodal-based discontinuous Galerkin solver, as was explained in Section 3.1. Therefore, we have to focus on the only thing that is common for every DGSEM solver: the solution is computed at the Gauss nodes inside each element. Hence, the state is simply defined as the values of the variable of interest, $q$, at the Gauss nodes, as represented in Figure 5(a). The size of the resulting state is $p+1$ (the same as the number of Gauss nodes) and is different depending on the polynomial order. This approach considers a different state in each element of the computational mesh; that is, the RL agent will handle each element independently, leading to a mesh-independent strategy. However, the main issue behind this definition of the state is that, in a general problem, $q$ may take any value, since $q\in\mathbb{R}$. This would lead to an unbounded and continuous state space, which does not match neither the MDP framework nor the value iteration algorithm, as they both deal with a finite state space. Hence, the possible values of the variable of interest $q$ must be discretized, so we can define an appropriate finite space state $\mathcal{S}$.
The proposed methodology takes the values of $q$ at every node within an element and normalizes the values from $-1$ to $1$. In this way, for each element there will always be one node with a value of $-1$ and another with a value of $1$. If it is the case that the value of $q$ is exactly the same at every node (below a very small tolerance), then the value at each node is set to 00 (see Section 3.5 for additional implementation details). Then, the state is discretized in $N_{l}$ levels. The higher the value of $N_{l}$, the more accurate the final agent will be, but the slower the training phase. After some trial and error, we discovered that a value of $N_{l}=11$ provides a good trade-off.
@@ 3.4.3 Reward definition
The definition of the reward will be influenced by the framework used to implement the p-adaptation strategy. In general, the p-adaptation process in a high-order solver is highly time-consuming. Therefore, we cannot afford to adapt the mesh every iteration. Furthermore, for numerical stability reasons, each time the mesh is adapted, only one action is taken in each element; that is, the polynomial order can be increased or decreased by one unit (see Section 3.5 for additional details regarding the implementation). These details should be considered when defining an appropriate definition of the reward function.
Additionally, we want the agent to find the optimal polynomial order for each element of the computational mesh, but it is not relevant how it reaches that polynomial, and therefore the reward must not depend on the action. However, it is important to decide whether the reward should depend on the current state $s$, or on the next state $s^{\prime}$. Both approaches are possible and valid, but the meaning is slightly different in each case. The key behind this decision is based on the knowledge that the trained agent will rely on the value function $v(s)$ to follow the optimal policy, which is simply to select the action that leads to a state with a higher $v(s^{\prime})$ (see the previous Section 3.3). Furthermore, as it is stated in eq. (2), each value function is directly related to the reward $r(s,a,s^{\prime})$, and hence there is a direct correlation between the selected reward and the final policy.
@@
Within this framework, if we choose the second reward option, $r=r(s^{\prime})$, the agent will try to find a state from which it is possible to reach a highly beneficial future state. However, if we can only perform one action at a time, the agent will not be able to reach the desired state until we can adapt a second time, at least. By the time we adapt again (several iterations later), the flow field might have changed, and that desirable future state might not be beneficial anymore. In contrast, if we define the reward as $r=r(s)$, the agent will take the action that leads to a highly beneficial state right now, even if the future is uncertain. In this way, at least, we can ensure that the chosen action is the best possible one right now. This is the reason why we have selected this approach when defining the reward function.
@@
where $\sigma$ is the standard deviation and $\alpha$ is a control parameter. The most important contribution to the reward function is the exponential part, which measures the accuracy (or the error). This part will drop the reward to zero if the $rmse$ is large compared to the standard deviation, $\sigma$. The second contribution is the cost, and it will only be relevant when the $rmse$ is small enough. This part will decrease the reward when a high polynomial order is used, as it will require a high computational cost. Therefore, the reward function prioritizes the reduction of the error; but if the error is acceptable (based on the value of $\sigma$), then the polynomial degree is reduced as much as possible (without significantly increasing the error).
On the one hand, the value of the parameter $\alpha$ provides a way to measure the trade-off between accuracy and computational cost. When $\alpha$ is small, the accuracy is more important than cost and vice versa. On the other hand, the value of the parameter $\sigma$ provides an estimate of the error that must be achieved in one element before the cost is taken into account. Therefore, the smaller the value of $\sigma$, the more important is the precision over the cost. Based on preliminary tests (not shown here), we have selected $\alpha=0.9$ and $\sigma=0.05$, which provides an appropriate balance between efficiency and accuracy of the p-adaptation strategy.